FTIR SPECTRA OF AZIDE ION IN REVERSE MICELLES

\underline{Q. ZHONG}, D. A. STEINHURST, J. C. OWRUTSKY, Naval Research Laboratory, Code 6111, Chemistry Division, Washington DC 20375-5342.

Reverse micelles (RM) are nanosize water droplets solubilized in bulk nonpolar organic solvent by ambiphilic surfactants. Confinement effects and solvation properties can be investigated by changing the water content (ω=[H_{2}O]/[surfactant]). FTIR spectroscopy was used to study the antisymmetric stretching vibration band of azide ion (2048.5 cm^{−1} in bulk water) dissolved in nonionic nonylphenol polyoxyethylene RM as a function of ω. The azide band is red-shifted (compared to in bulk water) and tends toward the bulk value with increasing ω. The band position also depends on the charge of the surfactant. The shift with ω is compared with changes observed for other RM properties and is explained in terms of the water core micropolarity.